# Frequency-Dependent Squeezing for Gravitational-Wave Detection through Quantum Teleportation

### v2

@@ V Discussions
Our scheme attains frequency-dependent squeezing by replacing the external filter cavities, which would otherwise require kilometer-scale additional vacuum tunnel and suspension systems, with the main interferometer itself through the application of the quantum-teleportation technique. It offers a sensitivity advantage around the optical spring resonance, which is designed by the motivation to create a dip at very-low frequencies in the sensitivity. This benefits from the long effective length of the arm cavity as the filter cavities and contributes to expand the bandwidth of the whole detector to lower frequencies.
However, we also need to note two drawbacks of the QT squeezing: (1) a factor of 3 (4.8 dB) higher injected squeezing required to reach the same level of detected squeezing at the readout port as achieved by the baseline FDS. (2) threefold noise contributions from input and output losses, which limit the sensitivity across the entire frequency band (see the figure of noise budget in Appendix D).
In addition, it is essential to highlight the technical flexibility of our scheme. The configuration illustrated in Fig. 2 allows for seamless transitions between detuned and tuned configurations. This transition can be achieved by turning off the pumping laser beam for Victor and adjusting the pumping frequency of the other beam, along with the macroscopic lengths of the arm and signal extraction cavity,effectively reverting back to the EPR squeezing in [28]. Furthermore, in the case of detuned configurations, QT squeezing provides the capability to address variations in the SEC detuning by optimizing the parameters of pumping frequencies and macroscopic length tuning. This means we do not need to replace the input mirror of filter cavity itself corresponding to a change of the value of detuning.
@@ Acknowledgements.
Authors are grateful to S. Hild for stimulating discussions and insight into potential benefits our scheme might bring for Einstein Telescope. Research by Y. N. is supported by JSPS Grant-in-Aid for JSPS Fellows Grant Number 23KJ0787. T. Z. acknowledges the support of the Institute for Gravitational Wave Astronomy at the University of Birmingham. S. D. is grateful for support to the Faculty of Science and Engineering of Maastricht University and to the European Research Council (within the frames of the ERC-2020-AdG agreement No. 101019978).
@@ Appendix A Derivation of Wiener filters
Appendix B Mathematical description of Bell measurement
@@
$L(t)$ $=De^{i\{\omega_{0}+(\Delta_{a}+\Delta_{v})/2\}}+\mathrm{h.c.},$ (15)
@@
where $-\xi_{\mathrm{d}}+\frac{\pi}{2}+\arg D$ and $\zeta_{V}=\xi_{\mathrm{d}}+\frac{\pi}{2}+\arg D$ (see also Eqs. (9)-(12) in [46]). In the frequency domain, one obtains:
@@
Appendix C Interferometer response for idlers
@@ C.1 Parameter searching
$\Delta_{v,a}$ $=\frac{c}{2L_{\mathrm{SEC}}}\Big{[}\arccos\left(\frac{T_{\mathrm{SEM}}\frac{\gamma_{\mathrm{1arm}}}{\gamma_{1,2}}-1-R_{\mathrm{SEM}}}{2\sqrt{R_{\mathrm{SEM}}}}\right)+(2n_{v,a}-1)\pi+\phi_{0}\Big{]},$ (26)
@@
We have four tunable parameters: macroscopic arm and SEC length, $\delta L_{\mathrm{arm}}=q\lambda/2$ and $\delta L_{\mathrm{SEC}}=p\lambda/2$ where $q$ and $p$ are integer values, and $n_{v,a}$ meaning the number of the SEC spectral ranges contained in $\Delta_{v,a}$. In Table 2, we provide the optimal length and frequency tuning parameters. The adjustments required for both the arm and SEC are approximately 1.6 and 2.4 cm, respectively. Detuning frequencies of Victor and Alice are approximately $\Delta_{a}\sim 32.9$ MHz and $\Delta_{v}\sim-789$ MHz. Fig 7 illustrates a comparison of phase rotations, highlighting the ponderomotive squeezing and the approximate rotations achieved through the combination of the two idlers.
@@
$T_{\mathrm{SEM}}$ SEM power transmittance 90 %
@@
$\delta L_{\mathrm{arm}}$ Arm length tuning $10497\lambda$
$\delta L_{\mathrm{SEC}}$ SEC length tuning $15790\lambda$
@@
$\Delta_{v}$ Detuning of Victor $655\ \mathrm{kHz}+1059\ \mathrm{FSR}_{\mathrm{SEC}}$
@@
C.2 Effect of arm cavity and SEC loss
@@
Appendix D Noise budget
@@
Fig. 8 displays the quantum noise budget for the QT squeezing with squeezing level of -15 dB, alongside the corresponding ETLF results for conventional squeezing. Vacuum fields are introduced into the beam path by each loss source, and their total contribution is obtained by integrating the fields from the same loss point. Four optical losses have been accounted for, including (1) 4% injection loss that considers losses in the OPA cavity and Faraday isolator; (2) A loss of 45ppm assumed for the round-trip of light in the arm, as discussed in Section C.2. (3) A loss of 1000ppm is assumed for the signal extraction mirror, the central beam splitter, and the imperfections of the Michelson interferometer, included in the overall SEC loss. (4) A 3% loss in readout due to the combined losses in the Faraday isolator, output mode cleaner, and inefficiency of the photo-detector. Losses from any source result in uncorrelated vacuum noise in the beam path that reduces the level of squeezing and correlation between EPR-entangled photons.

### v3

@@ V Discussions
Our scheme attains frequency-dependent squeezing by replacing the arbitrary number of external filter cavities, which would otherwise require kilometer-scale additional vacuum tunnel and suspension systems, with the main interferometer itself through the application of the quantum-teleportation technique. When applying to ETLF, we revealed that it offers a sensitivity advantage around the optical spring resonance which is the primary objective of detuning. This benefits from the long effective length of the arm cavity as the filter cavities and contributes to expand the bandwidth of the whole detector to lower frequencies.
However, we also need to note two drawbacks of the QT squeezing: (1) a factor of 3 (4.8 dB) higher injected squeezing required to reach the same level of detected squeezing at the readout port as achieved by the baseline FDS. (2) threefold noise contributions from input and output losses, which limit the sensitivity across the entire frequency band (see the figure of noise budget in Appendix E). More generally, if $N$ states participates in the teleportation, the eventual squeezing level is degraded and noise contributions from input and output loss are increased both by a factor of $N$ in power.
In addition, it is essential to highlight the technical flexibility of our scheme. The configuration illustrated in Fig. 2 allows for seamless transitions between detuned and tuned configurations. This transition can be achieved by turning off the pumping laser beam for Victor and adjusting the pumping frequency of the other beam, along with the macroscopic lengths of the arm and signal extraction cavity,effectively reverting back to the EPR squeezing in [33]. Furthermore, QT squeezing provides the capability to address variations in the SEC detuning by optimizing the parameters of pumping frequencies and macroscopic length tuning. This means we do not need to replace the input mirror of the filter cavity itself to adjust the filter-cavity bandwidth, which might otherwise be required due to practical factors, such as changes of the main laser power or ice formation on the cryogenic mirrors.
@@ Acknowledgements.
Authors are grateful to S. Hild for stimulating discussions and insight into potential benefits our scheme might bring for Einstein Telescope, and the ANU CGA squeezer group for fruitful discussion about experimental realization. Research by Y. N. is supported by JSPS Grant-in-Aid for JSPS Fellows Grant Number 23KJ0787. T. Z. acknowledges the support of the Institute for Gravitational Wave Astronomy at the University of Birmingham. S. D. is grateful for support to the Faculty of Science and Engineering of Maastricht University and to the European Research Council (within the frames of the ERC-2020-AdG agreement No. 101019978).
@@ Appendix A Derivation of Wiener filters
Appendix B Generalization to N phase rotations
Fig. 6 illustrates quantum teleportation squeezing to generate arbitrary number of phase rotations. There are two possible approaches: connecting EPR squeezing and QT squeezing to the QT squeezing protocol as Fig. 6(a) and (b). The first protocol prepares three phase rotations, while the second prepares four. Inductively, one can achieve an arbitrary number of pre-filtering phases by preparing multiple entanglement states and iterating Bell measurements. When one prepares $N$ phase rotations, Wiener filters, $\mathbf{g}=\begin{pmatrix}g_{1}&g_{2}&\cdots&g_{N}\end{pmatrix}^{\top}$, can be expressed as follows:
$\mathbf{g}=\mathbf{S}_{\alpha\alpha}^{-1}\mathbf{S}_{B\alpha},$ (13)
where $\mathbf{S}_{\alpha}$ and $\mathbf{S}_{B\alpha}$ are the cross-spectral density matrix and vector, respectively, defined as:
$\mathbf{S}_{\alpha\alpha}=\begin{pmatrix}S_{\alpha_{1}\alpha_{1}}&S_{\alpha_{1}\alpha_{2}}&\cdots&S_{\alpha_{1}\alpha_{N}}\\ S_{\alpha_{2}\alpha_{1}}&S_{\alpha_{2}\alpha_{2}}&\cdots&S_{\alpha_{2}\alpha_{N}}\\ \vdots&\vdots&\ddots&\vdots\\ S_{\alpha_{N}\alpha_{1}}&S_{\alpha_{N}\alpha_{2}}&\cdots&S_{\alpha_{N}\alpha_{N}}\end{pmatrix},$ (14)
and
$\mathbf{S}_{B\alpha}=\begin{pmatrix}S_{B\alpha_{1}}\\ S_{B\alpha_{2}}\\ \vdots\\ S_{B\alpha_{N}}\end{pmatrix}.$ (15)
Figure 6: Schematics of the generalized quantum teleportation squeezing with (a) three equivalent-filter cavities and (b) four filter cavities. Here, $\mathbb{U}_{1}$ represents the ponderomotive squeezing and other operations represent phase rotations, and $\pmb{\hat{v}}_{\mathrm{sqz}}$ is a squeezed vacuum.
Appendix C Mathematical description of Bell measurement
@@
$L(t)$ $=De^{i\{\omega_{0}+(\Delta_{a}+\Delta_{v})/2\}t}+\mathrm{h.c.},$ (18)
@@
where $\zeta_{A}=-\xi_{\mathrm{d}}+\frac{\pi}{2}+\arg D$ and $\zeta_{V}=\xi_{\mathrm{d}}+\frac{\pi}{2}+\arg D$ (see also Eqs. (9)-(12) in [52]). In the frequency domain, one obtains:
@@
Appendix D Interferometer response for idlers
@@ D.1 Parameter searching
$\Delta_{v,a}$ $=\frac{c}{2L_{\mathrm{SEC}}}\Big{[}\arccos\left(\frac{T_{\mathrm{SEM}}\frac{\gamma_{\mathrm{1arm}}}{\gamma_{1,2}}-1-R_{\mathrm{SEM}}}{2\sqrt{R_{\mathrm{SEM}}}}\right)$
$\quad+(2n_{v,a}-1)\pi+\phi_{0}\Big{]},$ (29)
@@
We have four tunable parameters: macroscopic arm and SEC length, $\delta L_{\mathrm{arm}}=q\lambda/2$ and $\delta L_{\mathrm{SEC}}=p\lambda/2$ where $q$ and $p$ are integer values, and $n_{v,a}$ meaning the number of the SEC spectral ranges contained in $\Delta_{v,a}$. In Table 2, we provide the optimal length and frequency tuning parameters. The adjustments required for both the arm and SEC are approximately 4.6 and -2.1 cm, respectively. Detuning frequencies of Victor and Alice are approximately $\Delta_{a}\sim 319$ MHz and $\Delta_{v}\sim 962$ MHz and the corresponding OPA cavity length is $\sim 94$ cm. Fig 8 illustrates a comparison of phase rotations, highlighting the ponderomotive squeezing and the approximate rotations achieved through the combination of the two idlers.
@@
$T_{\mathrm{SEM}}$ SEM power transmittance 20 %
@@
$\delta L_{\mathrm{arm}}$ Arm length tuning $30000\lambda$
$\delta L_{\mathrm{SEC}}$ SEC length tuning $14030\lambda$
@@
$\Delta_{v}$ Detuning of Victor $656\ \mathrm{kHz}+642\ \mathrm{FSR}_{\mathrm{SEC}}$
@@
D.2 Effect of arm cavity and SEC loss
@@
Appendix E Noise budget
@@
Fig. 9 displays the quantum noise budget for the QT squeezing with squeezing level of -15 dB, alongside the corresponding ETLF results for conventional squeezing. Vacuum fields are introduced into the beam path by each loss source, and their total contribution is obtained by integrating the fields from the same loss point. Four optical losses have been accounted for, including (1) 4% injection loss that considers losses in the OPA cavity and Faraday isolator; (2) A loss of 45ppm assumed for the round-trip of light in the arm, as discussed in Section D.2. (3) A loss of 1000ppm is assumed for the signal extraction mirror, the central beam splitter, and the imperfections of the Michelson interferometer, included in the overall SEC loss. (4) A 3% loss in readout due to the combined losses in the Faraday isolator, output mode cleaner, and inefficiency of the photo-detector. Losses from any source result in uncorrelated vacuum noise in the beam path that reduces the level of squeezing and correlation between EPR-entangled photons.